A 0.0600 kg ice cube at 30.0°C is placed in 0.537 kg of 35.0°C water in a very well insulated container. What is the final temperature? The latent heat of fusion of water is 79.8 kcal/kg, the specific heat of ice is 0.50 kcal/(kg · °C), and the specific heat of water is 1.00 kcal/(kg · °C).
Answer #1

T 6 Kg, T-300C 2 mw = 0.537 Kg 235 L 79.8Kallkg,1= 0.5 keallrg Cw=Ikal lkge final temperature, Mw CWT2-T)=m (0-Ti) tm + MrCw